OpenStudy (anonymous):

how to solve f(x)=2x^2-10 find f(5)

hartnn (hartnn):

to get f(5) just plug in x=5 in f(x) !

hartnn (hartnn):

so, what you get 2x^2-10 =2(5)^2-10 = ... ?

OpenStudy (anonymous):

would it be 90

hartnn (hartnn):

nopes, 5^2 = 25 2*25 = 50 50 -10 =40 it would be 40 :)

OpenStudy (anonymous):

ok thx for the help

hartnn (hartnn):

welcome ^_^